OpenStudy (anonymous):

The double integral of f=(x^2+y^2), x from -1 to 2 and y from 0 to sqrt(4-x^2). I need to do this with integration in polar coordinates but I have no idea how to express the cutoff at x=-1. Help!

OpenStudy (anonymous):

draw it.

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

split into two integrals here.

OpenStudy (anonymous):

\[ x=-1 \implies r\cos\theta = -1 \implies r=-\sec\theta \]

OpenStudy (anonymous):

Is it absolutely necessary to split this in to 2 integrals? If so then I know what I'm doing wrong. Is there a way to solve it using a single double integral?

OpenStudy (anonymous):

You must split it up.

OpenStudy (anonymous):

I see. Thanks for your help!

OpenStudy (anonymous):

You have two \(r(\theta)\) function and I see no way of combining them nicely.